OpenStudy (anonymous):

What is the product of 4x – 3 and x + 8?

OpenStudy (anonymous):

Just need to foil them. Or distribute.

OpenStudy (anonymous):

Are you familiar with the process?

OpenStudy (anonymous):

\[4x ^{2}+32x-3x-24\]

OpenStudy (anonymous):

yes i was jus confused on tht .. thanks polpak

OpenStudy (anonymous):

anwer\[4x ^{2}-29x+24\] becomes

OpenStudy (anonymous):

confused about what?

OpenStudy (anonymous):

if i needed to do foil

OpenStudy (anonymous):

ah.. Yep.

OpenStudy (anonymous):

@ mathfrick .. it would actually be -24 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

foil tends to be extra baggage to remember; i think. just multiply them together: 4x –3 x +8 ------------- 32x -24 4x^2 -3x ------------- 4x^2+29x -24

OpenStudy (anonymous):

yeah -24 sorry 4 that mis understanding

OpenStudy (anonymous):

I agree. Foil is memory clutter. I personally prefer to use the same strategy I use on monomials: \[(4x-3)(x+8) = 4x(x+8) - 3(x+8) = 4x^2 + 32x - 3x - 24 =\ ...\]

OpenStudy (anonymous):

Plus this method works with any number of terms.

OpenStudy (anonymous):

thank u so much guys!